OpenStudy (bohotness):

9. If the radius of the circle above is 4 cm, what is the circumference of the circle in terms of ? A. 4 cm B. 8 cm C. 12 cm D. 16 cm

OpenStudy (bohotness):

@satellite73

OpenStudy (bohotness):

@hhelpplzzzz

OpenStudy (rational):

circumference = \(\large 2\times \pi \times r\)

OpenStudy (bohotness):



OpenStudy (bohotness):

@rational